OpenStudy (anonymous):

Solve for n: p = m + n / -5

OpenStudy (compassionate):

Hello, and WELCOME TO OPENSTUDY! \[ p = m + \frac{ n }{ -5 }\] IS that the equation

OpenStudy (anonymous):

No, it is \[p = (m + n) / (-5) \] In other words, the sum of m + n is being divided by -5, and the answer to that equals p.

OpenStudy (anonymous):

n = -5p - 5m

OpenStudy (anonymous):

Thanks.

OpenStudy (texaschic101):

no....I think it is p = (m + n)/-5 -- multiply the equation by -5) -5p = m + n -- subtract m from both sides) -5p - m = n

OpenStudy (anonymous):

Question, why would you multiply by -5 and why would you subtract m from both sides?

OpenStudy (texaschic101):

to get n by itself

OpenStudy (texaschic101):

you multiply by -5 to get rid of the fraction. subtract m from both sides to cancel the m's out on the right side and because you want n by itself.

OpenStudy (anonymous):

Yeah, I think your answer is right, I just did the work and your method made more sense.

OpenStudy (texaschic101):

lol....luv the pic......your wecome